# Phi-3 Technical Report: A Highly Capable Language Model Locally on Your Phone

### v3

@@ 6.1 Technical Specifications
The Phi-3-Vision (4.2B parameters) is a multimodal model designed to process an image and a textual prompt as inputs, and subsequently generate textual outputs. This model is composed of two primary components: an image encoder, i.e., CLIP ViT-L/14 [RKH+21] and a transformer decoder, i.e., phi-3-mini-128K-instruct. The visual tokens, once extracted by the image encoder, are then combined with text tokens in an interleaved way (no particular order for image and text tokens). To accommodate high-resolution images and various aspect ratios, a dynamic cropping strategy [DZZ+24b] is utilized to split the input image into a 2d array of blocks, where the tokens of the blocks are concatenated to represent the whole image.
@@
The Phi-3-Vision model undergoes a pre-training phase using a diverse dataset, which consists of a combination of interleaved image-text documents (e.g.,  [LST+24]), image-text pairs from FLD-5B  [XWX+24], synthetic data derived from Optical Character Recognition (OCR) of PDF files, datasets for chart/table comprehension, and text-only data. The objective of predicting the next token is employed specifically on text tokens, while any loss associated with image tokens is disregarded during this phase. The pre-training process involves a total of $0.5T$ tokens that encompass both visual and text elements. During the pre-training phase, the maximum image resolution is capped at $1344\times 1344$ as the majority of the training images are smaller than this resolution.
@@
The Phi-3-Vision model contains two post-training stages: supervised finetuning (SFT) and direct preference optimization (DPO). For SFT, we leveraged text SFT dataset, public multimodal instruct tuning datasets along with large-scale multimodal instruct tuning datasets that we built ourselves, covering diverse domains and tasks such as general natural image understanding, chart/table/diagram understanding/reasoning, PowerPoint understanding, and model safety. The multimodal SFT data has about a total of 15B tokens. For DPO we mainly use a text DPO dataset and a relatively smaller-scale multimodal DPO dataset. For these two stages, we jointly train multimodal tasks and text-only tasks so that the model can achieve multi-modal reasoning while maintaining language capabilities as much as possible.
@@ 6.2 Academic benchmarks
We report in Table 2 the evaluation results of Phi-3-Vision on nine open-source academic benchmarks. These benchmarks evaluate reasoning and perceptual capabilities on visual and text inputs and can be grouped in three categories: Science, Charts, and Generic knowledge. We compare Phi-3-Vision with the following baselines: MM1-3B-Chat [MGF+24], MM1-7B-Chat [MGF+24], Llava-1.6 Vicuna 7B [LLLL23], Llava-1.6 Llama3-8B [LLL+24], Qwen-VL-Chat [BBY+23], Claude 3 Haiku [Ant24], Gemini 1.0 Pro V [TAB+23], and GPT-4V-Turbo. Our performance quality assessment setup used the same evaluation pipeline for all the baselines to ensure a fair comparison, with the exception of MM1-3B-Chat. We just copied and pasted their published numbers since the model is not publicly available.
@@
Our evaluation setup aimed to mimic scenarios where regular users interact with a multi-modal model, i.e., users who are not experts in prompt engineering or know special techniques that can improve performance. For this reason, we adopted the evaluation setting used in Llava-1.5 [LLLL23]. In this setup, the prompts include instructions to select a single letter corresponding to an answer from a list of given options, or answer with a single word or phrase. In our prompts, we did not use specific tokens for multiple-choice questions. Moreover, we did not scale or pre-process any image in our benchmarking system. We placed the images as the first item in the prompts, except on the MMMU dataset where the prompts interleave the images anywhere in the question or the answers. Lastly, our evaluation setup only considered a 0-shot format. Because of these evaluation parameters, our reported numbers can differ from the published numbers of the considered baselines.
@@
Phi-3-Vision 4.2b MM1-3B-Chat 3.6b [MGF+24] MM1-7B-Chat 7.6b [MGF+24] LLaVA-1.6 Vicuna-7b [LLLL23] LLaVA-Next LLama3-8b [LLL+24] Qwen-VL-Chat 9.6b [BBY+23] Claude 3 haiku  [Ant24] Gemini 1.0 Pro V  [TAB+23] GPT-4V-Turbo turbo-2024-04-09 MMMU (val) [YNZ+23] 40.4 33.9 37.0 34.2 36.4 39.0 40.7 42.0 55.5 ScienceQA (test) [LMX+22] 90.8 69.4 72.6 70.6 73.7 67.2 72.0 79.7 75.7 MathVista (testmini) [LBX+24] 44.5 32.0 35.9 31.5 34.8 29.4 33.2 35.0 47.5 Inter-GPS (test) [LGJ+21] 38.1 - - 20.5 24.6 22.3 32.1 28.6 41.0 MMBench (dev-en) [LDZ+24] 80.5 75.9 79.0 76.3 79.4 75.8 62.4 80.0 86.1 POPE (test) [LDZ+23] 85.8 87.4 86.6 87.2 87.0 82.6 74.4 84.2 83.7 AI2D (test) [KSK+16] 76.7 - - 63.1 66.9 59.8 60.3 62.8 74.7 ChartQA (test) [MLT+22] 81.4 - - 55.0 65.8 50.9 59.3 58.0 62.3 TextVQA (test) [SNS+19] 70.9 71.9 72.8 64.6 55.7 59.4 62.7 64.7 68.1
@@ 6.3 Safety
To ensure the integration of Phi-3-Vision aligns with Microsoft’s Responsible AI (RAI) principles, we involved safety post-training in both Supervised Fine-Tuning (SFT) stage and Direct Preference Optimization (DPO) stage. In creating the safety training datasets, we utilized not only the text-only RAI datasets, but also a variety of in-house Multi-Modal (MM) RAI datasets that cover various harm categories identified in both public and internal MM RAI benchmarks. For the purpose of RAI evaluation, we performed a rigorous quantitative assessment on both public and internal benchmarks, this was done in conjunction with a human evaluation conducted by Microsoft’s internal red team.
Phi-3-Vision 3.8b+0.3b Phi-3-Vision w/o safety 3.8b+0.3b Llava-1.6 Vicuna 7b+0.3b Qwen-VL-Chat 7.7b+1.9b GPT4-V N/A Internal (private) 8.30 7.06 5.44 7.27 8.55 RTVLM (public) 4.64 3.56 3.86 4.78 6.81 VLGuard (public) 9.12 4.66 5.62 8.33 8.90
@@
In Table 3, we present the evaluation outcomes of Phi-3-Vision on three MM RAI benchmarks: one internal and two public benchmarks (specifically, RTVLM [LLY+24] and VLGuard [ZBY+24]). We juxtapose these results with those of other open-source models such as Llava-1.5 [LLLL23], Llava-1.6 [LLL+24], Qwen-VL-Chat [BBY+23], and GPT4-V[Ope23]. The results clearly indicate that safety post-training notably enhances the RAI performance of Phi-3-Vision across all RAI benchmarks. In Figure 7, we further breakdown the performance across different RAI categories of the VLGuard and Internal benchmarks, demonstrating that safety post-training can aid Phi-3-Vision in improving RAI performance in nearly all categories.
Figure 7: Comparison of categorized RAI performance of Phi-3-Vision with and without the safety post-training on the VLGuard (left) and Internal (right) benchmark, respectively. It clearly indicates that safety post-training can enhance the RAI performance across nearly all the RAI categories.
@@ 6.4 Weakness
Regarding the multi-modal LLM capabilities of our Phi-3-Vision, it performs admirably across various fields. However, we have identified certain limitations, particularly with questions necessitating high-level reasoning abilities. Additionally, the model has been observed to occasionally generate ungrounded outputs, making it potentially unreliable in sensitive areas, such as finance. To mitigate these issues, we will incorporate more reasoning-focused and hallucination-related DPO data into post-training in the future.
From a responsible AI standpoint, whilst safety post-training has made significant strides, our Phi-3-Vision occasionally fails to refrain from answering harmful or sensitive inquiries. Examples of such occasions include deciphering particular types of captcha and describing scam images containing disinformation or hallucination. We find that this issue partly arises from the capabilities, such as OCR, acquired during the training process with normal instruct tuning datasets, which can be regarded as the trade-off between helpfulness and harmlessness. Moving forward, we need to further explore this area to achieve a better balance.
@@ Appendix B Authors (alphabetical)
Marah Abdin Jamie Huynh Olli Saarikivi
Sam Ade Jacobs Dan Iter Amin Saied
Ammar Ahmad Awan Russell J. Hewett Adil Salim
Jyoti Aneja Mojan Javaheripi Michael Santacroce
Hany Awadalla Xin Jin Shital Shah
Ahmed Awadallah Nikos Karampatziakis Ning Shang
Nguyen Bach Piero Kauffmann Hiteshi Sharma
Amit Bahree Mahoud Khademi Swadheen Shukla
Arash Bakhtiari Dongwoo Kim Xia Song
Jianmin Bao Lev Kurilenko Yin Tat Lee
@@
Alon Benhaim Yunsheng Li Xin Wang
Misha Bilenko Chen Liang Guanhua Wang
Johan Bjorck Lars Liden Lijuan Wang
Sébastien Bubeck Eric Lin Chunyu Wang
Caio César Teodoro Mendes Zeqi Lin Yu Wang
Martin Cai Weishung Liu Rachel Ward
Qin Cai Mengchen Liu Wen Wen
Vishrav Chaudhary Ce Liu Philipp Witte
Weizhu Chen Chong Luo Haiping Wu
Yi-Ling Chen Li Lyna Zhang Michael Wyatt
Yen-Chun Chen Piyush Madan Bin Xiao
Dongdong Chen David Majercak Jiahang Xu
Dong Chen Matt Mazzola Can Xu
Parul Chopra Arindam Mitra Weijian Xu
Xiyang Dai Hardik Modi Sonali Yadav
Gustavo de Rosa Anh Nguyen Fan Yang
Allie Del Giorno Brandon Norick Ziyi Yang
Matthew Dixon Barun Patra Jianwei Yang
Ronen Eldan Daniel Perez-Becker Yifan Yang
Victor Fragoso Thomas Portet Donghan Yu
Mei Gao Reid Pryzant Lu Yuan
Min Gao Heyang Qin Chengruidong Zhang
Jianfeng Gao James R. Lee Cyril Zhang
Abhishek Goswami Marko Radmilac Jianwen Zhang
Suriya Gunasekar Corby Rosset Yi Zhang
Emman Haider Sambudha Roy Yunan Zhang
Junheng Hao Olatunji Ruwase Xiren Zhou
@@ Appendix C Acknowledgements
We would like to thank Zhuohan Li, Simon Mo, and Kaichao You from UC Berkeley for sharing their insights on the vLLM kernel.

### v4

@@ 7.1 Technical Specifications
The Phi-3.5-Vision (4.2B parameters) is a multimodal model designed to process an image/multi-image and a textual prompt as inputs, and subsequently generate textual outputs. This model is composed of two primary components: an image encoder, i.e., CLIP ViT-L/14 [RKH+21] and a transformer decoder, i.e., phi-3.5-mini. The visual tokens, once extracted by the image encoder, are then combined with text tokens in an interleaved way (no particular order for image and text tokens). To accommodate high-resolution images and various aspect ratios, a dynamic cropping strategy [DZZ+24b] is utilized to split the input image into a 2d array of blocks, where the tokens of the blocks are concatenated to represent the whole image. For multi-image input, we simply concatenated tokens from each images together.
@@
The Phi-3.5-Vision model undergoes a pre-training phase using a diverse dataset, which consists of a combination of interleaved image-text documents (e.g.,  [LST+24]), image-text pairs from FLD-5B  [XWX+24], synthetic data derived from Optical Character Recognition (OCR) of PDF files, datasets for chart/table comprehension, and text-only data. The objective of predicting the next token is employed specifically on text tokens, while any loss associated with image tokens is disregarded during this phase. The pre-training process involves a total of $0.5T$ tokens that encompass both visual and text elements. During the pre-training phase, the maximum image resolution is capped at $1344\times 1344$ as the majority of the training images are smaller than this resolution.
@@
The Phi-3.5-Vision model contains two post-training stages: supervised finetuning (SFT) and direct preference optimization (DPO). For SFT, we leveraged text SFT dataset, public multimodal instruct tuning datasets along with large-scale multimodal instruct tuning datasets that we built ourselves, covering diverse domains and tasks such as general natural image understanding, chart/table/diagram understanding/reasoning, PowerPoint understanding, multi-image comparison, video summarization and model safety. The multimodal SFT data has about a total of 33B tokens. For DPO we mainly use a text DPO dataset and a relatively smaller-scale multimodal DPO dataset. For these two stages, we jointly train multimodal tasks and text-only tasks so that the model can achieve multi-modal reasoning while maintaining language capabilities as much as possible.
@@ 7.2.1 Single-image Benchmarks
We report in Table 5 the evaluation results of Phi-3.5-Vision on nine open-source academic benchmarks. These benchmarks evaluate reasoning and perceptual capabilities on visual and text inputs and can be grouped in three categories: Science, Charts, and Generic knowledge. We compare Phi-3.5-Vision with the following baselines: MM1-3B-Chat [MGF+24], MM1-7B-Chat [MGF+24], Llava-1.6 Vicuna 7B [LLLL23], Llava-1.6 Llama3-8B [LLL+24], Qwen-VL-Chat [BBY+23], Claude 3 Haiku [Ant24], Gemini 1.0 Pro V [TAB+23], and GPT-4O. Our performance quality assessment setup used the same evaluation pipeline for all the baselines to ensure a fair comparison, with the exception of MM1-3B-Chat. We just copied and pasted their published numbers since the model is not publicly available.
@@
Our evaluation setup aimed to mimic scenarios where regular users interact with a multi-modal model, i.e., users who are not experts in prompt engineering or know special techniques that can improve performance. For this reason, we adopted the evaluation setting used in Llava-1.5 [LLLL23]. In this setup, the prompts include instructions to select a single letter corresponding to an answer from a list of given options, or answer with a single word or phrase. In our prompts, we did not use specific tokens for multiple-choice questions. Moreover, we did not scale or pre-process any image in our benchmarking system. We placed the images as the first item in the prompts, except on the MMMU dataset where the prompts interleave the images anywhere in the question or the answers. Lastly, our evaluation setup only considered a 0-shot format. Because of these evaluation parameters, our reported numbers can differ from the published numbers of the considered baselines. As we can seen, our Phi-3.5-Vision achieves super competitive results on all benchmarks and outperform other competitor models on most benchmarks while being smaller.
7.2.2 Multi-image Benchmarks
We report in Table 6 the evaluation results of Phi-3.5-Vision on one latest academic multi-image benchmark and one video benchmark. These benchmarks evaluate perceptual capabilities on multiple image/frames and text covering a wide range of general scenarios (e.g., Art and Style recognition, Forensic detection, and video understanding). We compare Phi-3.5-Vision with the following baseline methods: Llava Interleave-Qwen 7B [LZZ+24], InternVL2 4B and 8B [CWT+24], Gemini 1.5 Flash [TAB+23], GPT-4o-mini, Claude 3.5 Sonnet [Ant24], Gemini 1.5 Pro [TAB+23], and GPT-4O. Line in the single-frame evaluation case, our performance quality assessment setup used the same evaluation pipeline for all the baselines to ensure a fair comparison.
Our evaluation setup for multi-image also followed the Llava setup where prompts include instructions to select a single letter corresponding to an answer from a list of given options, or answer with a single word or phrase. Moreover, we did not use specific tokens for multiple-choice questions and we did not scale or pre-process any image in our benchmarking system. For most of the benchmarks, we placed the images as the first item in the prompts.
The evaluation pipelines for BLINK and VideoMME benchmarks differ from those published. In the case of BLINK, we do not use ChatGPT as the final answer selection mechanism. Instead, we instruct the evaluated model to select one answer directly from the given choices. The reason is that in this manner we ensure that the mistakes or successes come solely by the evaluated model. For the VideoMME benchmark, we extracted 16 frames from the video by sampling frames at a given rate that ensures a uniform time coverage of the entire video. We used 16 frames since this is the maximum number of images a prompt can contain for Azure OpenAI models. Unlike the proposed evaluation in VideoMME that uses the maximum number of frames a model can accept, we always pass the same amount of frames across all the considered model baselines. In this way we ensure the evaluations are fair since all the models receive the exact same input information (i.e., the prompt and set of images). As shown in Table 6, our Phi-3.5-Vision performs very competitively or outperforms baseline models under the similar model size in multi-image understanding scenarios as well.
@@ 7.2.2 Multi-image Benchmarks
Phi-3.5-Vision 4.2b MM1-3B-Chat 3.6b [MGF+24] MM1-7B-Chat 7.6b [MGF+24] LLaVA-1.6 Vicuna-7b [LLLL23] LLaVA-Next LLama3-8b [LLL+24] Qwen-VL-Chat 9.6b [BBY+23] Claude 3 haiku  [Ant24] Gemini 1.0 Pro V  [TAB+23] GPT-4O 2024-05-13 MMMU (val) [YNZ+23] 43.0 33.9 37.0 34.2 36.4 39.0 40.7 42.0 61.8 ScienceQA (test) [LMX+22] 91.3 69.4 72.6 70.6 73.7 67.2 72.0 79.7 88.5 MathVista (testmini) [LBX+24] 43.9 32.0 35.9 31.5 34.8 29.4 33.2 35.0 54.4 Inter-GPS (test) [LGJ+21] 36.3 - - 20.5 24.6 22.3 32.1 28.6 46.9 MMBench (dev-en) [LDZ+24] 81.9 75.9 79.0 76.3 79.4 75.8 62.4 80.0 88.4 POPE (test) [LDZ+23] 86.1 87.4 86.6 87.2 87.0 82.6 74.4 84.2 87.0 AI2D (test) [KSK+16] 78.1 - - 63.1 66.9 59.8 60.3 62.8 82.8 ChartQA (test) [MLT+22] 81.8 - - 55.0 65.8 50.9 59.3 58.0 64.0 TextVQA (test) [SNS+19] 72.0 71.9 72.8 64.6 55.7 59.4 62.7 64.7 75.6
@@
Phi-3.5-Vision 4.2b Llava-interleave Qwen 7b [LZZ+24] InternVL2 4b [CWT+24] InternVL2 8b [CWT+24] Gemini 1.5 Flash [TAB+23] GPT4O mini 2024-07-18 Claude 3.5 Sonnet  [Ant24] Gemini 1.5 Pro  [TAB+23] GPT-4O 2024-05-13 BLINK (val) [FHL+24] 57.0 53.1 45.9 45.4 45.8 51.9 56.5 61.0 63.2 VideoMME (test) [FDL+24] 50.8 50.2 49.9 52.6 62.3 61.2 55.9 62.6 68.4
Table 6: Comparison results on public multi-image/video MLLM benchmarks. All the reported numbers are produced with the exact same pipeline to ensure that the numbers are comparable.
@@ 7.3 Safety
To ensure the integration of Phi-3.5-Vision aligns with Microsoft’s Responsible AI (RAI) principles, we involved safety post-training in both Supervised Fine-Tuning (SFT) stage and Direct Preference Optimization (DPO) stage. In creating the safety training datasets, we utilized not only the text-only RAI datasets, but also a variety of in-house Multi-Modal (MM) RAI datasets that cover various harm categories identified in both public and internal MM RAI benchmarks. For the purpose of RAI evaluation, we performed a rigorous quantitative assessment on both public and internal benchmarks, this was done in conjunction with a human evaluation conducted by Microsoft’s internal red team.
Phi-3.5-Vision 3.8b+0.3b Phi-3.5-Vision w/o safety 3.8b+0.3b Llava-1.6 Vicuna 7b+0.3b Qwen-VL-Chat 7.7b+1.9b GPT4-V N/A Internal (private) 8.16 7.06 5.44 7.27 8.55 RTVLM (public) 5.44 3.56 3.86 4.78 6.81 VLGuard (public) 9.10 4.66 5.62 8.33 8.90
@@
In Table 7, we present the evaluation outcomes of Phi-3.5-Vision on three MM RAI benchmarks: one internal and two public benchmarks (specifically, RTVLM [LLY+24] and VLGuard [ZBY+24]). We juxtapose these results with those of other open-source models such as Llava-1.5 [LLLL23], Llava-1.6 [LLL+24], Qwen-VL-Chat [BBY+23], and GPT4-V[Ope23]. The results clearly indicate that safety post-training notably enhances the RAI performance of Phi-3.5-Vision across all RAI benchmarks. In Figure 8, we further breakdown the performance across different RAI categories of the VLGuard and Internal benchmarks, demonstrating that safety post-training can aid Phi-3.5-Vision in improving RAI performance in nearly all categories.
Figure 8: Comparison of categorized RAI performance of Phi-3.5-Vision with and without the safety post-training on the VLGuard (left) and Internal (right) benchmark, respectively. It clearly indicates that safety post-training can enhance the RAI performance across nearly all the RAI categories.
@@ 7.4 Weakness
Regarding the multi-modal LLM capabilities of our Phi-3.5-Vision, it performs admirably across various fields. However, we have identified certain limitations, particularly with questions necessitating high-level reasoning abilities. Additionally, the model has been observed to occasionally generate ungrounded outputs, making it potentially unreliable in sensitive areas, such as finance. To mitigate these issues, we will incorporate more reasoning-focused and hallucination-related DPO data into post-training in the future.
From a responsible AI standpoint, whilst safety post-training has made significant strides, our Phi-3.5-Vision occasionally fails to refrain from answering harmful or sensitive inquiries. Examples of such occasions include deciphering particular types of captcha and describing scam images containing disinformation or hallucination. We find that this issue partly arises from the capabilities, such as OCR, acquired during the training process with normal instruct tuning datasets, which can be regarded as the trade-off between helpfulness and harmlessness. Moving forward, we need to further explore this area to achieve a better balance.
@@ Appendix B Authors (alphabetical)
Marah Abdin Xin Jin Adil Salim
Jyoti Aneja Nikos Karampatziakis Michael Santacroce
Hany Awadalla Piero Kauffmann Shital Shah
Ahmed Awadallah Mahoud Khademi Ning Shang
Ammar Ahmad Awan Dongwoo Kim Hiteshi Sharma
Nguyen Bach Young Jin Kim Yelong Shen
Amit Bahree Lev Kurilenko Swadheen Shukla
Arash Bakhtiari James R. Lee Xia Song
Jianmin Bao Yin Tat Lee Masahiro Tanaka
@@
Alon Benhaim Yunsheng Li Praneetha Vaddamanu
Misha Bilenko Chen Liang Chunyu Wang
Johan Bjorck Lars Liden Guanhua Wang
Sébastien Bubeck Xihui Lin Lijuan Wang
Martin Cai Zeqi Lin Shuohang Wang
Qin Cai Ce Liu Xin Wang
Vishrav Chaudhary Liyuan Liu Yu Wang
Dong Chen Mengchen Liu Rachel Ward
Dongdong Chen Weishung Liu Wen Wen
Weizhu Chen Xiaodong Liu Philipp Witte
Yen-Chun Chen Chong Luo Haiping Wu
Yi-Ling Chen Piyush Madan Xiaoxia Wu
Hao Cheng Ali Mahmoudzadeh Michael Wyatt
Parul Chopra David Majercak Bin Xiao
Xiyang Dai Matt Mazzola Can Xu
Matthew Dixon Caio César Teodoro Mendes Jiahang Xu
Ronen Eldan Arindam Mitra Weijian Xu
Victor Fragoso Hardik Modi Jilong Xue
Jianfeng Gao Anh Nguyen Sonali Yadav
Mei Gao Brandon Norick Fan Yang
Min Gao Barun Patra Jianwei Yang
Amit Garg Daniel Perez-Becker Yifan Yang
Allie Del Giorno Thomas Portet Ziyi Yang
Abhishek Goswami Reid Pryzant Donghan Yu
Suriya Gunasekar Heyang Qin Lu Yuan
Emman Haider Marko Radmilac Chenruidong Zhang
Junheng Hao Liliang Ren Cyril Zhang
Russell J. Hewett Gustavo de Rosa Jianwen Zhang
Wenxiang Hu Corby Rosset Li Lyna Zhang
Jamie Huynh Sambudha Roy Yi Zhang
Dan Iter Olatunji Ruwase Yue Zhang
Sam Ade Jacobs Olli Saarikivi Yunan Zhang
Mojan Javaheripi Amin Saied Xiren Zhou
@@ Appendix C Acknowledgements
We would like to thank Zhuohan Li, Simon Mo from UC Berkeley and Kaichao You from Tsinghua University for sharing their insights on the vLLM kernel.